Circular No. 8393

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 Q1 (TUCKER) Roy A. Tucker, Tucson, AZ, reports his discovery of a comet with a

coma diameter of ∼ 50′′ and a 70′′ tail in p.a. ∼ 230^{o} (discovery observation below) on unfiltered CCD images taken with a 0.35-m reflector. Following posting on the ‘NEO Confirmation Page’, the object was also noted to be cometary on CCD images taken by J. Foster (University Hills, Los Angeles, CA, 0.32-m f/5.6 reflector; fan tail in p.a. 235^{o} on Aug. 24.4 UT) and by J. E. McGaha (Tucson, AZ, 0.36-m f/10 reflector; faint 26′′ coma with a very faint fan-shaped tail 30′′ long in p.a. 250^{o} on Aug. 25.35). Images by Tucker on Aug. 25.5 show a 30′′ coma and a tail ∼ 40′′ long in p.a. 255^{o}.

2004 UT  α_{2000}  δ_{2000}  Mag.
Aug. 23.45566  2^{h}42^{m}20.^{s}76  +2^{o}08′54′′.4  14.6

The available astrometry and the following preliminary parabolic orbital elements appear on MPEC 2004-Q43 (ephemeris: H = 10.0, n = 3).

T = 2004 Dec. 10.500 TT  ω = 35.^{o}562
Ω = 21.955 }2000.0
q = 2.00387 AU  i = 56.610
2004 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  Mag.
Aug. 23 2^{h}42.^{m}36 + 2^{o}01.^{′}0 1.841 2.394 110.^{o}7 23.^{o}3 14.2
28 2 41.82 + 3 30.2 1.748 2.362 115.2 22.8 14.0
Sept. 2 2 40.41 + 5 06.6 1.658 2.332 120.0 22.0 13.9
7 2 38.01 + 6 51.0 1.572 2.303 124.9 21.0 13.7
12 2 34.48 + 8 44.3 1.491 2.274 130.2 19.8 13.5
17 2 29.69 +10 47.2 1.415 2.247 135.7 18.2 13.4
22 2 23.53 +12 59.9 1.346 2.221 141.4 16.4 13.2
27 2 15.88 +15 22.0 1.285 2.196 147.3 14.3 13.1
Oct. 2 2 06.68 +17 52.4 1.232 2.172 153.1 12.0 13.0

SUPERNOVAE 2004ch AND 2004do M. Salvo and B. Schmidt, Australian National University (ANU), re-

port on preliminary reduction of spectra taken with the ANU 2.3-m telescope (+ Double-Beam spectrograph) at Siding Spring: SN 2004do (cf. IAUC 8382), Aug. 19.37 UT, type-Ia supernova, ∼ 10 days past maximum light; SN 2003ch (cf. IAUC 8353), Aug. 20.58, type-II supernova, a couple of months past maximum light.

2004 August 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green